# Object-Centric Representation Learning for Enhanced 3D Semantic Scene Graph Prediction

### v2

@@ 1 Introduction
Figure 1: (a) VL-SAT wang2023vl embeds object features non-discriminatively, leading to low-confidence predictions and frequent object misclassifications, which degrade relationship accuracy. In contrast, (b) our method embeds object features in a more discriminative manner, yielding high confidence scores and more accurate object classifications. Consequently, relationship predictions are significantly improved, resulting in a more coherent and semantically accurate scene graph.
Recent research in 3D Semantic Scene Graph (3DSSG) Prediction has significantly enhanced the semantic understanding of 3D environments. By abstracting raw point cloud data into structured semantic graphs that capture objects and their interrelationships, 3DSSG facilitates critical tasks such as robot navigation gu2024conceptgraphs; yin2024sg; gomez2020hybrid, object manipulation dhamo2021graph, and VR/AR interactions tahara2020retargetable. This technology has become essential for semantic-level 3D understanding in various applications sarkar2023sgaligner; agia2022taskography; nuthalapati2021lightweight; azuma_2022_CVPR; etesam20223dvqa by reducing ambiguity and facilitating intuitive human-machine interaction.
A variety of deep learning approaches have been proposed for accurate 3DSSG prediction armeni20193d; kim20193; gay2019visual; rosinol2021kimera; wald2022learning; wu2023incremental; zhang2021sggpoint; fanfan20223d. SGPN wald2020learning pioneered the point cloud-based relationship prediction framework and released the 3DSSG dataset, which we utilize in this study. SGFN wu2021scenegraphfusion introduced a feature-wise attention layer for more accurate inference and proposed an incremental update methodology. Most recently, VL-SAT wang2023vl addressed the problem of long-tailed distribution utilizing visual and textual information. Despite these advances, we identify two critical limitations in current approaches.
@@
Lack of elaborating relationship information. Prior works woo2018linknet; tang2019learning; yang2018graph; sharifzadeh2021classification; li2017vip; liao2019natural; yu2017visual; zhou2019visual often fail to effectively integrate object information when constructing relationship features for prediction. Methods like SGFN wu2021scenegraphfusion, VL-SAT wang2023vl, and 3D-VLAP wang2024weakly rely solely on geometric relationship information between objects as edge features, neglecting object semantic characteristics. Conversely, approaches like Zhang et al. zhang2021knowledge and SGPN wald2020learning use PointNet-based features from scene-level point clouds containing both objects, introducing excessive background information that can degrade prediction accuracy.
@@ 2 Observations
SGPN wald2020learning 8% 12% 18%
SGFN wu2021scenegraphfusion 8% 12% 20%
VL-SAT wang2023vl 8% 13% 19%
@@
Eq. (1) makes explicit that sharper object posteriors $P(o\mid\mathbf{z})$—i.e., more discriminative embeddings—yield lower-entropy mixture and thus higher confidence in predicate prediction. In contrast, ambiguous object embeddings result in broader $P(o\mid\mathbf{z})$ distributions, diluting the contribution of correct object labels and increasing the likelihood of predicate errors. Similarly, prior studies in vision tasks—such as marino2016more; fang2017object; deng2014large—have demonstrated that incorporating prior object knowledge aids relation prediction. Notably, chen2019knowledge leveraged the statistical co-occurrence of objects and explicitly regularized the semantic relationship prediction to address the uneven distribution issue. Compared to other works, we elaborated these assumptions with discriminative object features. For a more detailed comparison and analysis of our method, please refer to Section B of the Appendix.
@@ 3 Proposed Methods
Our goal is to predict a 3D semantic scene graph represented as a directed graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ where $\mathcal{V}$ and $\mathcal{E}$ denote the sets of nodes and edges, respectively. The input consists of a 3D point cloud $\mathbf{P}\in\mathbb{R}^{N\times 3}$, containing $N$ points, a set of class-agnostic instance masks $\mathcal{M}=\{M_{k}\}_{k=1}^{K}$, which associate the point cloud with $K$ semantic instances. Also, 3DSSG takes the scan data of 3RScan  wald2019rio which contains RGB-D sequence when constructing the 3D point cloud data. Here, we can extract a collection of RGB images $\mathcal{I}_{k}=\{I_{1},I_{2},...,I_{n_{k}}\}$ for each object instance in $\hat{o}_{k}\in\mathcal{V}$ following the procedure in VL-SAT wang2023vl. Each object instance $\hat{o}_{i}$ represents the nodes in scene graph with ground-truth label $o_{i}\in\mathcal{O}$, while each edge $e_{ij}\in\mathcal{E}$ encodes the predicates in a triplet $\langle$ subject, predicate, object $\rangle$, where the head node $\hat{o}_{i}$ serves as the subject and the tail node $\hat{o}_{j}$ as the object. Each $\hat{o}_{i}$ corresponds to one of $N_{obj}$ semantic object classes, while each edge $e_{ij}$ can contain multiple predicate labels from $N_{pred}$ semantic relation classes such as ‘standing on’, ‘hanging in’, and others.
@@ 3.1 Object Feature Learning
Discriminative object features with image/text modals. As noted in Section 2, our objective is to pre-train an encoder that extracts discriminative object features, thereby enabling more confident object classification. To this end, we employ a contrastive pre-training scheme inspired by CrossPoint afham2022crosspoint and $\text{CLIP}^{2}$ zeng2023clip2, leveraging the text descriptions and 2D RGB images available for each 3D object instance. For an instance $\hat{o}_{i}$ with RGB views $\mathcal{I}_{i}=\{I_{n}\}_{n=1}^{n_{i}}$, we cast the task as a contrastive learning problem with multiple positive samples per anchor. Features from both modalities are obtained with a pretrained CLIP radford2021learning ViT-B/32 encoder $f_{\theta_{c}}(\cdot)$. Specifically, each image is encoded as $\mathbf{z}_{\text{image}}^{i}\in\mathcal{Z}_{I}^{i}=\{f_{\theta_{c}}(I):I\in\mathcal{I}_{i}\}$, while the textual description $T_{\hat{o}_{i}}$ is encoded as $\textbf{z}_{\text{text}}^{i}=f_{\theta_{c}}(T_{\hat{o}_{i}})$. The text prompt follows the template “A point cloud of {object}.”
Our objective is to maximize the similarity between the object embedding vector $\mathbf{z}^{t}$ and its corresponding image and text features $\mathbf{z}^{i}_{\text{image}},\textbf{z}_{\text{text}}^{i}$, while minimizing similarity to other embeddings in the mini-batch. Following previous works yeh2022decoupled; koch2024lang3dsg; zeng2023clip2, we adopt a supervised contrastive framework to efficiently utilize label data, assuming that objects with the same label share semantic characteristics across their point clouds and associated images. However, unlike text prompts which provide a single representation per object, multiple images can correspond to a single object instance. To handle this asymmetry, we separate the contrastive losses for visual and textual modality. Let $\mathcal{D}_{B}=\{(\mathbf{z}^{t}_{i},o_{i})\}_{i=1}^{B}$ denote the batch of $B$ point cloud features and their corresponding ground-truth labels. For each index $i\in\mathcal{B}=\{1,...,B\}$, we define the set of positive indices as $\mathcal{P}(i)=\{p\in\mathcal{B}:o_{p}=o_{i}\}$, and the set of negatives as $\mathcal{N}(i)=\mathcal{B}/\mathcal{P}(i)$. The visual contrastive loss is defined as:
@@
Unlike prior works afham2022crosspoint; zeng2023clip2, we do not include the positive samples in the denominator, thereby promoting a more robust and discriminative feature space. This design promotes the pretrained encoder to extract more discriminative features, which reduces conditional uncertainty $H(o\mid\mathbf{z})$ and leads to lower predicate prediction errors (please refer to the Appendix Section B for more details).
@@ 3.2 Relationship Feature Learning
After the pretraining stage, we train the relationship feature encoder and the GNNs for scene graph prediction. Unlike prior approaches, our relationship feature encoder integrates both object-level semantic features and their geometric relationships. Following the design of geometric descriptors used in prior works wang2023vl; wu2021scenegraphfusion, we adopt their formulation to represent the spatial configuration between object pairs. The geometric descriptor between objects is defined as:
@@ 4 Experiments
Datasets and task descriptions. We evaluate our approach on the 3DSSG dataset wald2020learning, a semantically enriched extension of 3RScan wald2019rio designed for 3D semantic scene graph prediction ***We used the 3DSSG and 3RScan dataset wald2020learning under the permission of its creators and authors. We contacted the authors by email and google form.. This dataset consists of 1,553 real-world indoor scenes, annotated with 160 object categories and 26 predicate types, covering a wide range of household environments. We follow the standard train/validation split defined in the original benchmark. For details of the evaluation metrics, please refer to Section D of the Appendix.
@@ 4.1 Comparison with Other Works
We compare our method with the state-of-the-art approach VL-SAT wang2023vl, as well as several existing baselines, including SGPN wald2020learning and SGFN wu2021scenegraphfusion, using their publicly available implementations or reported results. Our evaluation includes both quantitative performance metrics and qualitative analysis, including t-SNE visualizations of object feature spaces and 3D scene graph visualizations. Additional experimental results are provided in Section D of the Appendix.
@@
SGPN wald2020learning 49.46 73.99 86.92 94.76 85.38 88.59
SGFN wu2021scenegraphfusion 53.36 76.88 89.00 97.71 88.59 91.14
VL-SAT wang2023vl 55.93 78.06 89.81 98.46 89.35 92.20
@@
Quantitative comparison. Tables 2 and 3 show that our model achieves a new state-of-the-art performance on the 3DSSG benchmark across all evaluation metrics. The most significant improvements are observed in object classification. As reported in Table 2, our method improves object classification accuracy by 2–4% compared to VL-SAT wang2023vl. This enhanced object representation contributes directly to better relationship classification, supporting our main assumption. Our method also consistently outperforms baselines in both SGCls and PredCls tasks, as reported in Table 3. Regardless of whether graph constraints are applied, we observe performance gains of 1–4% in both settings. These consistent improvements demonstrate not only the effectiveness of our approach over existing methods, but also provide empirical support for our assumption: enhancing object discrimination leads to more precise relationship prediction. The improvements in PredCls, where object labels are given and only relational reasoning is required, further substantiate this effect.
Model SGCls (w/ GC) PredCls (w/ GC) SGCls (w/o GC) PredCls (w/o GC) R@20 R@50 R@100 R@20 R@50 R@100 R@20 R@50 R@100 R@20 R@50 R@100 SGPN wald2020learning 27.0 28.8 29.0 51.9 58.0 58.5 28.2 32.6 35.3 54.5 70.1 82.4 Zhang et al. zhang2021knowledge 28.5 30.0 30.1 59.3 65.0 65.3 29.8 34.3 37.0 62.2 78.4 88.3 SGFN wu2021scenegraphfusion 29.5 31.2 31.2 65.9 78.8 79.6 31.9 39.3 45.0 68.9 82.8 91.2 VL-SAT wang2023vl 32.0 33.5 33.7 67.8 79.9 80.8 33.8 41.3 47.0 70.5 85.0 92.5 Ours 36.1 37.7 37.8 70.2 82.0 82.6 38.1 46.1 52.5 73.3 87.8 94.6
@@
Qualitative comparison. Fig. 5 presents qualitative comparisons of predicted scene graphs between our method and VL-SAT wang2023vl. VL-SAT frequently misclassifies visually similar but semantically distinct objects such as cabinet and chair, or stool and garbage bin, which leads to erroneous relationship predictions. These misclassifications result in hallucinated relationships that do not exist in the ground truth—for instance, predicting $\langle$ desk, standing on, floor $\rangle$ or a spurious relation between shelf and cabinet. While such relationships might appear plausible given the visual similarity, they ultimately stem from incorrect object recognition, which propagates errors to the relational reasoning module. In contrast, our method correctly identifies object categories, thereby facilitating accurate and consistent relationship prediction. These examples highlight the critical role of object classification in downstream relation inference and reinforce the importance of robust object-level representations for reliable 3D scene graph construction.
Figure 6: The t-SNE visualization of the object latent space of the VL-SAT wang2023vl (a) and our model (b).
@@ 4.2 Discussions
Plug-in our object encoder to other models. To verify that our performance gains considerably stem from the object encoder, we integrated our pretrained encoder into three representative baselines: SGPN wald2020learning, SGFN wu2021scenegraphfusion, and VL-SAT wang2023vl, while keeping all other components unchanged. The results in Table 4 show consistent improvements across all frameworks. For SGPN wald2020learning, object recall increases at most 7% and other metrics also increased significantly. Most notably, even the state-of-the-art VL-SAT wang2023vl benefits from our object encoder, where overall metrics are improved. These consistent enhancements across different architectures confirm that our pretraining strategy yields more discriminative and transferable object embeddings that directly benefit relationship prediction.

### v3

@@ 1 Introduction
Figure 1: (a) VL-SAT Wang et al. (2023) embeds object features non-discriminatively, leading to low-confidence predictions and frequent object misclassifications, which degrade relationship accuracy. In contrast, (b) our method embeds object features in a more discriminative manner, yielding high confidence scores and more accurate object classifications. Consequently, relationship predictions are significantly improved, resulting in a more coherent and semantically accurate scene graph.
Recent research in 3D Semantic Scene Graph (3DSSG) Prediction has significantly enhanced the semantic understanding of 3D environments. By abstracting raw point cloud data into structured semantic graphs that capture objects and their interrelationships, 3DSSG facilitates critical tasks such as robot navigation Gu et al. (2024); Yin et al. (2024); Gomez et al. (2020), object manipulation Dhamo et al. (2021), and VR/AR interactions Tahara et al. (2020). This technology has become essential for semantic-level 3D understanding in various applications Sarkar et al. (2023); Agia et al. (2022); Nuthalapati et al. (2021); Azuma et al. (2022); Etesam et al. (2022) by reducing ambiguity and facilitating intuitive human-machine interaction.
A variety of deep learning approaches have been proposed for accurate 3DSSG prediction Armeni et al. (2019); Kim et al. (2019); Gay et al. (2019); Rosinol et al. (2021); Wald et al. (2022); Wu et al. (2023); Zhang et al. (2021a); Fanfan et al. (2022). SGPN Wald et al. (2020) pioneered the point cloud-based relationship prediction framework and released the 3DSSG dataset, which we utilize in this study. SGFN Wu et al. (2021) introduced a feature-wise attention layer for more accurate inference and proposed an incremental update methodology. Most recently, VL-SAT Wang et al. (2023) addressed the problem of long-tailed distribution utilizing visual and textual information. Despite these advances, we identify two critical limitations in current approaches.
@@
Lack of elaborating relationship information. Prior works Woo et al. (2018); Tang et al. (2019); Yang et al. (2018); Sharifzadeh et al. (2021); Li et al. (2017); Liao et al. (2019); Yu et al. (2017); Zhou et al. (2019) often fail to effectively integrate object information when constructing relationship features for prediction. Methods like SGFN Wu et al. (2021), VL-SAT Wang et al. (2023), and 3D-VLAP Wang et al. (2024) rely solely on geometric relationship information between objects as edge features, neglecting object semantic characteristics. Conversely, approaches like Zhang et al. Zhang et al. (2021b) and SGPN Wald et al. (2020) use PointNet-based features from scene-level point clouds containing both objects, introducing excessive background information that can degrade prediction accuracy.
@@ 2 Observations
SGPN Wald et al. (2020) 8% 12% 18%
SGFN Wu et al. (2021) 8% 12% 20%
VL-SAT Wang et al. (2023) 8% 13% 19%
@@
Eq. (1) makes explicit that sharper object posteriors $P(o\mid\mathbf{z})$—i.e., more discriminative embeddings—yield lower-entropy mixture and thus higher confidence in predicate prediction. In contrast, ambiguous object embeddings result in broader $P(o\mid\mathbf{z})$ distributions, diluting the contribution of correct object labels and increasing the likelihood of predicate errors. Similarly, prior studies in vision tasks—such as Marino et al. (2016); Fang et al. (2017); Deng et al. (2014)—have demonstrated that incorporating prior object knowledge aids relation prediction. Notably, Chen et al. (2019) leveraged the statistical co-occurrence of objects and explicitly regularized the semantic relationship prediction to address the uneven distribution issue. Compared to other works, we elaborated these assumptions with discriminative object features. For a more detailed comparison and analysis of our method, please refer to Section B of the Appendix.
@@ 3 Proposed Methods
Our goal is to predict a 3D semantic scene graph represented as a directed graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ where $\mathcal{V}$ and $\mathcal{E}$ denote the sets of nodes and edges, respectively. The input consists of a 3D point cloud $\mathbf{P}\in\mathbb{R}^{N\times 3}$, containing $N$ points, a set of class-agnostic instance masks $\mathcal{M}=\{M_{k}\}_{k=1}^{K}$, which associate the point cloud with $K$ semantic instances. Also, 3DSSG takes the scan data of 3RScan  Wald et al. (2019) which contains RGB-D sequence when constructing the 3D point cloud data. Here, we can extract a collection of RGB images $\mathcal{I}_{k}=\{I_{1},I_{2},...,I_{n_{k}}\}$ for each object instance in $\hat{o}_{k}\in\mathcal{V}$ following the procedure in VL-SAT Wang et al. (2023). Each object instance $\hat{o}_{i}$ represents the nodes in scene graph with ground-truth label $o_{i}\in\mathcal{O}$, while each edge $e_{ij}\in\mathcal{E}$ encodes the predicates in a triplet $\langle$ subject, predicate, object $\rangle$, where the head node $\hat{o}_{i}$ serves as the subject and the tail node $\hat{o}_{j}$ as the object. Each $\hat{o}_{i}$ corresponds to one of $N_{obj}$ semantic object classes, while each edge $e_{ij}$ can contain multiple predicate labels from $N_{pred}$ semantic relation classes such as ‘standing on’, ‘hanging in’, and others.
@@ 3.1 Object Feature Learning
Discriminative object features with image/text modals. As noted in Section 2, our objective is to pre-train an encoder that extracts discriminative object features, thereby enabling more confident object classification. To this end, we employ a contrastive pre-training scheme inspired by CrossPoint Afham et al. (2022) and $\text{CLIP}^{2}$ Zeng et al. (2023), leveraging the text descriptions and 2D RGB images available for each 3D object instance. For an instance $\hat{o}_{i}$ with RGB views $\mathcal{I}_{i}=\{I_{n}\}_{n=1}^{n_{i}}$, we cast the task as a contrastive learning problem with multiple positive samples per anchor. Features from both modalities are obtained with a pretrained CLIP Radford et al. (2021) ViT-B/32 encoder $f_{\theta_{c}}(\cdot)$. Specifically, each image is encoded as $\mathbf{z}_{\text{image}}^{i}\in\mathcal{Z}_{I}^{i}=\{f_{\theta_{c}}(I):I\in\mathcal{I}_{i}\}$, while the textual description $T_{\hat{o}_{i}}$ is encoded as $\textbf{z}_{\text{text}}^{i}=f_{\theta_{c}}(T_{\hat{o}_{i}})$. The text prompt follows the template “A point cloud of {object}.”
Our objective is to maximize the similarity between the object embedding vector $\mathbf{z}^{t}$ and its corresponding image and text features $\mathbf{z}^{i}_{\text{image}},\textbf{z}_{\text{text}}^{i}$, while minimizing similarity to other embeddings in the mini-batch. Following previous works Yeh et al. (2022); Koch et al. (2024a); Zeng et al. (2023), we adopt a supervised contrastive framework to efficiently utilize label data, assuming that objects with the same label share semantic characteristics across their point clouds and associated images. However, unlike text prompts which provide a single representation per object, multiple images can correspond to a single object instance. To handle this asymmetry, we separate the contrastive losses for visual and textual modality. Let $\mathcal{D}_{B}=\{(\mathbf{z}^{t}_{i},o_{i})\}_{i=1}^{B}$ denote the batch of $B$ point cloud features and their corresponding ground-truth labels. For each index $i\in\mathcal{B}=\{1,...,B\}$, we define the set of positive indices as $\mathcal{P}(i)=\{p\in\mathcal{B}:o_{p}=o_{i}\}$, and the set of negatives as $\mathcal{N}(i)=\mathcal{B}/\mathcal{P}(i)$. The visual contrastive loss is defined as:
@@
Unlike prior works Afham et al. (2022); Zeng et al. (2023), we do not include the positive samples in the denominator, thereby promoting a more robust and discriminative feature space. This design promotes the pretrained encoder to extract more discriminative features, which reduces conditional uncertainty $H(o\mid\mathbf{z})$ and leads to lower predicate prediction errors (please refer to the Appendix Section B for more details).
@@ 3.2 Relationship Feature Learning
After the pretraining stage, we train the relationship feature encoder and the GNNs for scene graph prediction. Unlike prior approaches, our relationship feature encoder integrates both object-level semantic features and their geometric relationships. Following the design of geometric descriptors used in prior works Wang et al. (2023); Wu et al. (2021), we adopt their formulation to represent the spatial configuration between object pairs. The geometric descriptor between objects is defined as:
@@ 4 Experiments
Datasets and task descriptions. We evaluate our approach on the 3DSSG dataset Wald et al. (2020), a semantically enriched extension of 3RScan Wald et al. (2019) designed for 3D semantic scene graph prediction ***We used the 3DSSG and 3RScan dataset Wald et al. (2020) under the permission of its creators and authors. We contacted the authors by email and google form.. This dataset consists of 1,553 real-world indoor scenes, annotated with 160 object categories and 26 predicate types, covering a wide range of household environments. We follow the standard train/validation split defined in the original benchmark. For details of the evaluation metrics, please refer to Section D of the Appendix.
@@ 4.1 Comparison with Other Works
We compare our method with the state-of-the-art approach VL-SAT Wang et al. (2023), as well as several existing baselines, including SGPN Wald et al. (2020) and SGFN Wu et al. (2021), using their publicly available implementations or reported results. Our evaluation includes both quantitative performance metrics and qualitative analysis, including t-SNE visualizations of object feature spaces and 3D scene graph visualizations. Additional experimental results are provided in Section D of the Appendix.
@@
SGPN Wald et al. (2020) 49.46 73.99 86.92 94.76 85.38 88.59
SGFN Wu et al. (2021) 53.36 76.88 89.00 97.71 88.59 91.14
VL-SAT Wang et al. (2023) 55.93 78.06 89.81 98.46 89.35 92.20
@@
Quantitative comparison. Tables 2 and 3 show that our model achieves a new state-of-the-art performance on the 3DSSG benchmark across all evaluation metrics. The most significant improvements are observed in object classification. As reported in Table 2, our method improves object classification accuracy by 2–4% compared to VL-SAT Wang et al. (2023). This enhanced object representation contributes directly to better relationship classification, supporting our main assumption. Our method also consistently outperforms baselines in both SGCls and PredCls tasks, as reported in Table 3. Regardless of whether graph constraints are applied, we observe performance gains of 1–4% in both settings. These consistent improvements demonstrate not only the effectiveness of our approach over existing methods, but also provide empirical support for our assumption: enhancing object discrimination leads to more precise relationship prediction. The improvements in PredCls, where object labels are given and only relational reasoning is required, further substantiate this effect.
Model SGCls (w/ GC) PredCls (w/ GC) SGCls (w/o GC) PredCls (w/o GC) R@20 R@50 R@100 R@20 R@50 R@100 R@20 R@50 R@100 R@20 R@50 R@100 SGPN Wald et al. (2020) 27.0 28.8 29.0 51.9 58.0 58.5 28.2 32.6 35.3 54.5 70.1 82.4 Zhang et al. Zhang et al. (2021b) 28.5 30.0 30.1 59.3 65.0 65.3 29.8 34.3 37.0 62.2 78.4 88.3 SGFN Wu et al. (2021) 29.5 31.2 31.2 65.9 78.8 79.6 31.9 39.3 45.0 68.9 82.8 91.2 VL-SAT Wang et al. (2023) 32.0 33.5 33.7 67.8 79.9 80.8 33.8 41.3 47.0 70.5 85.0 92.5 Ours 36.1 37.7 37.8 70.2 82.0 82.6 38.1 46.1 52.5 73.3 87.8 94.6
@@
Qualitative comparison. Fig. 5 presents qualitative comparisons of predicted scene graphs between our method and VL-SAT Wang et al. (2023). VL-SAT frequently misclassifies visually similar but semantically distinct objects such as cabinet and chair, or stool and garbage bin, which leads to erroneous relationship predictions. These misclassifications result in hallucinated relationships that do not exist in the ground truth—for instance, predicting $\langle$ desk, standing on, floor $\rangle$ or a spurious relation between shelf and cabinet. While such relationships might appear plausible given the visual similarity, they ultimately stem from incorrect object recognition, which propagates errors to the relational reasoning module. In contrast, our method correctly identifies object categories, thereby facilitating accurate and consistent relationship prediction. These examples highlight the critical role of object classification in downstream relation inference and reinforce the importance of robust object-level representations for reliable 3D scene graph construction.
Figure 6: The t-SNE visualization of the object latent space of the VL-SAT Wang et al. (2023) (a) and our model (b).
@@ 4.2 Discussions
Plug-in our object encoder to other models. To verify that our performance gains considerably stem from the object encoder, we integrated our pretrained encoder into three representative baselines: SGPN Wald et al. (2020), SGFN Wu et al. (2021), and VL-SAT Wang et al. (2023), while keeping all other components unchanged. The results in Table 4 show consistent improvements across all frameworks. For SGPN Wald et al. (2020), object recall increases at most 7% and other metrics also increased significantly. Most notably, even the state-of-the-art VL-SAT Wang et al. (2023) benefits from our object encoder, where overall metrics are improved. These consistent enhancements across different architectures confirm that our pretraining strategy yields more discriminative and transferable object embeddings that directly benefit relationship prediction.